Hello, I'm reviewing for a final and need some help checking this problem:
Solve each equation:
n/3 - 2n/5 = -1
thanks!

Answer by RAY100(1637)

n/3 - 2n/5 = -1
.
multiply all terms by LCD =15
.
5n - 6n =-15
.
-1n = -15
.
n = 15,,,,,Answer
.
checking
.
15/3 -2*15 / 5 =-1
.
5 - 6 =-1 =-1 ,,,,,,ok